### v1

# Generalizing while preserving monotonicity in comparison-based preference learning models

Abstract
If you tell a learning model that you prefer an alternative $a$ over another alternative $b$, then you probably expect the model to be monotone, that is, the valuation of $a$ increases, and that of $b$ decreases. Yet, perhaps surprisingly, many widely deployed comparison-based preference learning models, including large language models, fail to have this guarantee. Until now, the only comparison-based preference learning algorithms that were proved to be monotone are the Generalized Bradley-Terry models DBLP:conf/aaai/FageotFHV24 . Yet, these models are unable to generalize to uncompared data. In this paper, we advance the understanding of the set of models with generalization ability that are monotone. Namely, we propose a new class of Linear Generalized Bradley-Terry models with Diffusion Priors, and identify sufficient conditions on alternatives’ embeddings that guarantee monotonicity. Our experiments show that this monotonicity is far from being a general guarantee, and that our new class of generalizing models improves accuracy, especially when the dataset is limited.
1 Introduction
Preference learning, sometimes known as alignment, has become central to machine learning. In particular, in recent years, there has been a growing interest to leverage comparative judgments to fine-tune AI models, using frameworks like Reinforcement Learning with Human Feedback (RLHF) DBLP:conf/nips/ChristianoLBMLA17 or Direct Preference Optimization (DPO) rafailovDirectPreferenceOptimization2024 in the context of language models, or linear models in the context of applications ranging from trolley dilemmas to food donation awadMoralMachineExperiment2018 ; DBLP:journals/pacmhci/LeeKKKYCSNLPP19 .These models are now deployed at scale.
Yet bizarre aspects of these preference learning algorithms are regularly observed. In particular, strikingly, DBLP:journals/corr/abs-2402-13228 ; DBLP:journals/corr/abs-2410-08847 ; pairwise-monotonicity showed that, when they are updated based on a comparative judgment saying that an alternative $a$ is preferable to $b$, the probability of generating item $a$ can in fact decrease. In fact, most deployed learning algorithms fail to guarantee monotonicity. A preferred alternative can be downgraded after it was said to be better than another alternative.
Perhaps surprisingly, the root cause of this lack of monotonicity is not the nonlinearity of the models. In fact, we can expose this “bug" with a very basic example. Consider a linear two-dimensional model with a parameter $\beta\in\mathbb{R}^{2}$ to be learned: the score of alternative $a$ and $b$ are $\beta^{\top}x_{a}$ and $\beta^{\top}x_{b}$, where $x_{a}$ and $x_{b}$ are (two-dimensional) vector embeddings. We are given a comparison that favors alternative $a$ over $b$, whose embeddings are $x_{a}=(1,0)$ and $x_{b}=(2,0)$. Since $x_{a1}<x_{b1}$, this comparison will push $\beta_{1}$ towards lower values. But since the score of $a$ according to the linear model is $\beta^{T}x_{a}=\beta_{1}$, this means that the score of $a$ will also decrease. Thus, including a comparison that favors $a$ over $b$ has decreased the score of $a$. This becomes all the more troubling when there exists an alternative $c$ with embedding $x_{c}=(0,1)$, whose score remains unchanged.
This example questions whether preference learning algorithms can be trusted. A user who witnesses a surprising evolution of alternatives’ scores as illustrated above might, understandably, prefer not to use such an algorithm. Worse yet, if they are nevertheless forced to use the algorithm, they could want to remove the data they previously provided, because the eventual learned model was deteriorated by their truthful data reporting. More generally, this may discourage users to report their preferences, and rather provide tactical preferences in the hope to steer the model towards their goal. This is reminiscent of tactical or “useful” strategies in voting systems. Now, some models, like the (Generalized) Bradley-Terry model, were proved to be monotone DBLP:conf/aaai/FageotFHV24 . But to the best of our knowledge, none of them enables score generalization to non-evaluated alternatives. Hence the following research problem.
Can a generalizing comparison-based preference learning algorithm guarantee monotonicity?
Contributions.
Our first contribution is to identify a large class of preference learning algorithms that leverage both comparisons between alternative pairs, and descriptive information (embeddings) on individual alternatives, which we call Linear Generalized Bradley-Terry with Diffusion Prior (Definition 3). This class extends the (Generalized) Bradley-Terry models DBLP:conf/aaai/FageotFHV24 by including a linear mapping of the embeddings—and priors on alternative similarities, thereby allowing preference generalization to yet uncompared alternatives.
As a second contribution, we provide conditions on the embeddings that guarantee that the learning algorithm behaves monotonically when new comparisons are provided. As discussed above, this property is highly desirable and yet hard to guarantee in practice. In particular, we propose a class of diffusion embeddings that guarantee monotony, and for which membership is easy-to-check. Interestingly, diffusion embeddings yield a very appealing interpretation as heat diffusion dynamics where comparisons play the role of heat pumps. A direct consequence is that categorical information (one-hot encoding embedding) yields a monotone learning algorithm. In particular, this class enables us to provide a positive answer to our research question.
Finally, we evaluate the statistical performance of our learning algorithms through numerical experiments on synthetic datasets. Our evaluations show that a linear model with good embeddings and diffusion priors outperforms the classical GBT model DBLP:conf/aaai/FageotFHV24 , in particular when the amount of data is limited.
Related works.
Learning preferences from comparisons has a long history, dating back to Thurstone thurstone27 , Zermelo zermelo1929berechnung , and Bradley and Terry bradley1952rank . To handle inconsistent judgments, such algorithms define a probabilistic model of how latent scoring of alternatives are transformed into noisy comparisons. Their approach was adapted by luce1959individual and plackett1975analysis to model the selection of one preferred alternative out of several proposed ones.
While the Bradley-Terry model models binary-valued comparisons, various authors have proposed extensions to real-valued comparisons, e.g., ranging the interval $[-1,1]$. Historically, this started with the modeling of draws in games like chess davidson1970extending . More recently,  DBLP:journals/corr/abs-1911-11658 proposed the platform Climpact where users are given pairs of activities, and are asked to evaluate the comparative pollutions of two activities. They then develop a model based on a quadratic error to turn the comparisons into evaluations of the amounts of pollution of the individual activities. Their model was then generalized by DBLP:conf/aaai/FageotFHV24 in a framework they call Generalized Bradley-Terry (GBT) to turn real-valued comparisons into scores. All these models however consider that each alternative’s score is an independent latent variable to be learned. Thus, they fail to generalize to non-evaluated alternatives.
Independent of user-provided comparisons, alternatives usually come with descriptive information. DBLP:journals/nca/MenkeM08 , csiszar2012algorithms and zhao2016deep proposed to model an alternative’s score as a parametrized function of a vector embedding of the description of the alternative. Recently, this trick has been widely used in the context of language models vaswani2017attention ; brown2020language , especially through algorithms like Reinforcement Learning with Human Feedback (RLHF)  DBLP:conf/nips/ChristianoLBMLA17 ; DBLP:conf/nips/StiennonO0ZLVRA20 , Direct Preference Optimization (DPO) dpo , or Generalized Preference Optimization (GPO) DBLP:conf/icml/TangGZCMRRVPP24 , to name a few. In this paper, we restrict ourselves to linear models, where an alternative’s score is assumed to be a linear function of their embedding. Such linear models of preferences trained from comparative judgments have previously been studied and used, e.g. by  DBLP:conf/ijcai/GuoTKOCCEDI18 ; DBLP:conf/aaai/NoothigattuGADR18 ; DBLP:journals/pacmhci/LeeKKKYCSNLPP19 .
The study of the mathematical guarantees of preference learning algorithms has only emerged recently. In particular, nonlinear models have been empirically shown to violate monotonicity properties DBLP:conf/nips/ChenMZ0ZRC24 ; DBLP:journals/corr/abs-2402-13228 ; DBLP:journals/corr/abs-2410-08847 . While pairwise-monotonicity proved that nonlinear models nevertheless provide a weak monotonicity guarantee they call local pairwise monotonicity, they also suggest that these models are unlikely to verify stronger forms of monotonicity. Conversely, and quite remarkably, DBLP:conf/aaai/FageotFHV24 proved that the GBT model guarantees monotonicity for all GBT root laws. Our model extends GBT in several ways.
Finally, in addition to AI model alignement, we note that learning preferences from comparisons with mathematical guarantees also contributes to developing more transparent social medial, as advocated recently by "prosocial media" weyl2025prosocialmedia . In particular, this work finds a direct application in the collaborative scoring of social media content hoang2021tournesolquestlargesecure ; hoang2024solidagomodularcollaborativescoring , and could extend the consensus-driven polling platform Pol.is small2021polis .
Paper structure.
The rest of the paper is organized as follows. Section 2 introduces the formalism, formally defines monotonicity, and recalls the GBT model. Section 3 defines the linear GBT model with diffusion prior, and states our main results. Section 4 provides the main lines of the proofs of the main results. Section 5 reports on our experiments. Section 6 concludes.
2 Monotonicity of Scoring Models
In this Section, we set notations, formalize the notion of monotonicity, and recall the Generalized Bradley-Terry model.
2.1 Notations and operations on datasets
Consider a set $[A]\triangleq\left\{1,2,\ldots,A\right\}$ of alternatives. The set $\mathcal{R}\subseteq\mathbb{R}$ denotes the set of admissible comparison values, which we assume to be symmetric around zero, i.e. $r\in\mathcal{R}\iff-r\in\mathcal{R}$. In the classical Bradley-Terry model, we have ={1,+$\mathcal{R}=\left\{-1,+1\right\}$. The generalized Bradley-Terry model allows a wider variety of possible comparison values, typically $\mathcal{R}=[-1,1]$. A comparison sample is defined as a triple $(a,b,r)$ where $a,b\in\mathcal{A}$ are two distinct alternatives, and $r\in\mathcal{R}$. We assume $(a,b,r)$ and $(b,a,-r)$ to be equivalent, which we write $(a,b,r)\simeq(b,a,-r)$. By also having $(a,b,r)\simeq(a,b,r)$ (and the relation false otherwise), we clearly constructed an equivalence relation. A dataset ${\bf D}$ is a list ${\bf D}:[N]\to[A]^{2}\times\mathcal{R}$ of comparison samples. We write $\mathcal{D}\triangleq\bigcup_{N\in\mathbb{N}}([A]^{2}\times\mathcal{R})^{N}$ for the set of datasets, and $\left\lvert{{\bf D}}\right\rvert$ the length of a dataset ${\bf D}\in\mathcal{D}$. We now define four parameterized operations $\mathcal{D}\to\mathcal{D}$ on datasets, and denote $\mathcal{O}(\mathcal{D})$ the set of compositions of the four operations.
Exchange.
For any $n\in\mathbb{N}$, $\operatorname*{\textsc{exchange}}_{n}({\bf D})$ is the dataset that only changes the $n$-th entry of the dataset (if it exists), by replacing $(a_{n},b_{n},r_{n})$ with $(a_{n},b_{n},-r_{n})$. All other entries are left unchanged. Assuming that preference learning algorithms should interpret these two comparison samples identically, this operation should not affect training.
Shuffle.
For any $N\in\mathbb{N}$ and any permutation $\sigma$ of $[N]$, $\operatorname*{\textsc{shuffle}}_{N,\sigma}$ acts on datasets by reordering its $N$ first elements (if $\left\lvert{{\bf D}}\right\rvert<N$, then ${\bf D}$ is left unchanged). Formally, assuming $\left\lvert{{\bf D}}\right\rvert\geq N$, for all $n\in[N]$ we have $\operatorname*{\textsc{shuffle}}_{N,\sigma}({\bf D})_{n}={\bf D}_{\sigma(n)}$. Again, if a preference learning algorithm is expected to be invariant to shuffling, then this operation should not affect training.
Append.
For any comparison sample $(a,b,r)$, we define $\operatorname*{\textsc{append}}_{a,b,r}({\bf D})$ the dataset that is obtained by appending $(a,b,r)$ to ${\bf D}$. Namely, we have $\left\lvert{\operatorname*{\textsc{append}}_{a,b,r}({\bf D})}\right\rvert=\left\lvert{{\bf D}}\right\rvert+1$ and $\operatorname*{\textsc{append}}_{a,b,r}({\bf D})_{\left\lvert{{\bf D}}\right\rvert+1}=(a,b,r)$, while other elements are the same as in ${\bf D}$. An append is said to infinitely favor $a^{\prime}$ over $b^{\prime}$ if it has parameters $(a,b,r)\simeq(a^{\prime},b^{\prime},\max\mathcal{R})$. Note that if $\mathcal{R}$ does not have a maximum, then no append infinitely favors $a^{\prime}$ over $b^{\prime}$.
Update.
For any $n\in\mathbb{N}$ and comparison $r\in\mathcal{R}$, $\operatorname*{\textsc{update}}_{n,r}({\bf D})$ replaces the comparison in the $n$-th entry of ${\bf D}$ with $r$ (if the $n$-th entry already existed). The update is said to favor $a^{\prime}$ over $b^{\prime}$ if one of the following occurs: (i) $(a_{n},b_{n})=(a^{\prime},b^{\prime})$ and $r\geq r_{n}$, or (ii) $(a_{n},b_{n})=(b^{\prime},a^{\prime})$ and $r\leq r_{n}$. In other words, it favors $a^{\prime}$ over $b^{\prime}$ if it acts on a comparison sample between $a^{\prime}$ and $b^{\prime}$, and modifies the comparison $r$ to further favor $a^{\prime}$.
2.2 Monotonicity
Definition 1 (Favoring $a$).
[…]
The relation $\preceq_{a}$ is clearly a preorder. Similarly, we define the preorder $\leq_{a}$ over $\mathbb{R}^{A}$ by $\theta\leq_{a}\theta^{\prime}$ if $\theta_{a}\leq\theta^{\prime}_{a}$ coordinate-wise. We can now formally define monotonicity.
Definition 2 (Monotonicity).
The preference learning algorithm Alg is monotone if, for every alternative $a\in[A]$, $\textsc{Alg}:(\mathcal{D},\preceq_{a})\to(\mathbb{R}^{A},\leq_{a})$ is monotone. More explicitly, for all $a$, if ${\bf D}\succeq_{a}{\bf D}^{\prime}$ then $\textsc{Alg}({\bf D})\geq_{a}\textsc{Alg}({\bf D}^{\prime})$.
Remark 1.
In the sequel, all preference learning algorithms that we will consider are neutral myerson2013fundamentals , i.e. they treat all alternatives symmetrically111 Formally, $(\sigma\cdot{\bf D})=\sigma\cdot\textsc{Alg}$ Alg for all permutations of $[A]$, with the action that applies pointwise to all apparitions of an alternative $a\in[A]$. . For such algorithms, the monotonicity for any single $a\in[A]$ clearly implies that for all $a\in[A]$.
2.3 (Generalized) Bradley-Terry and monotonicity
Here we recall the probabilistic model of GBT DBLP:conf/aaai/FageotFHV24 , slightly adapting it to our dataset formalism,.222 In DBLP:conf/aaai/FageotFHV24 , the authors consider datasets that contain at most one sample per pair of alternatives. Following Bradley and Terry bradley1952rank , GBT defines a probabilistic model of comparisons given scores. Specifically, given two alternatives $a$ and $b$ having scores $\theta_{a}$ and $\theta_{b}$, the probability of observing a value $r$ for the comparison of $a$ relative to $b$ is
$p(r|\theta_{ab})\propto f(r)\cdot\exp(r\cdot\theta_{ab}),$ (1)
where $\theta_{ab}=\theta_{a}-\theta_{b}$ is the score difference between $a$ and $b$, and $f$ is the root law, a probability distribution on $\mathcal{R}$ that describes comparisons when $a$ and $b$ have equal scores.
Given a dataset ${\bf D}=(a_{n},b_{n},r_{n})_{n\in[N]}$ of $N$ independent observations following (1), and assuming a gaussian prior with zero mean and $\sigma^{2}$ variance for each alternative score $\theta_{a}$, $a\in\mathcal{A}$, the standard Maximum A Posteriori methodology results in the GBT estimator:
$\textsc{GBT}_{f,\sigma}({\bf D})\triangleq\operatorname*{arg\,min}_{\theta\in\mathbb{R}^{\mathcal{A}}}\frac{1}{2\sigma^{2}}\sum_{a}\theta_{a}^{2}+\sum_{(a,b,r)\in{\bf D}}\Phi_{f}(\theta_{ab})-r\theta_{ab}.$ (2)
[…]
Theorem 2 DBLP:conf/aaai/FageotFHV24 shows that $\textsc{GBT}_{f,\sigma}$ is monotone, in a context where two elements can only be compared once. The forthcoming Theorem 3 shows this holds even when alternatives are repeated.
Proposition 1 (Slight generalization of DBLP:conf/aaai/FageotFHV24 ).
For any root law $f$, and $\sigma>0$, $\textsc{GBT}_{f,\sigma}$ is monotone.
Although well behaved in many aspects, the generalized Bradley-Terry model fails to perform generalization: an alternative $a$ that never appears in ${\bf D}$ will receive a nil score $\textsc{GBT}({\bf D})_{a}=0$. However, in practice, alternatives may (i) admit informative descriptions, and (ii) have known relationships. This should help guess the score of a yet uncompared alternative, based on the scoring of similar compared alternatives. We introduce such a learning algorithm in the next Section.
3 Linear GBT with Diffusion Prior
In this Section, we introduce a class of preference learning algorithms that incorporate both user comparisons and contextual information on the compared elements, and state our main result about their mathematical guarantees on monotonicity.
3.1 Learning with prior similarities
The $\textsc{GBT}_{\sigma,f}$ model does not include prior knowledge on the structure of the alternatives. For example, when alternatives represent videos on YouTube, the fact that two videos belong to the same channel cannot be represented in Equation (2). More generally, Equation (2) does not encode any prior similarities between alternatives. Consequently, if an alternative $a$ is never compared with any other, then, even if $a$ is similar to an alternative $b$ that has a large non-zero score, $a$ will still be assigned a zero score. In other words, Equation (2) does not allow us to generalize.
To address this issue, we generalize $\textsc{GBT}_{\sigma,f}$ (2) in two directions:
  1. 1.
    (Embeddings) We let $\theta=\theta(\beta)=x^{T}\beta$ be a (deterministic) linear function of a parameter $\beta\in\mathbb{R}^{D}$. The matrix $x\in\mathbb{R}^{D\times A}$ assigns an embedding $x_{a}\in\mathbb{R}^{D}$ to every alternative $a$. For instance, in the context of YouTube, $x$ could denote a one-hot encoding of the video-to-channel relation.
  2. 2.
    (Similarity) We consider a more general regularization term $\mathcal{R}(\beta)$ of the form
    $\mathcal{R}(\beta)=\frac{1}{2\sigma^{2}}\sum_{d}\beta_{d}^{2}+\frac{1}{2}\sum_{ab}\theta_{a}(\beta)L_{ab}\theta_{b}(\beta)$ (3)
    where $L$ is a Laplacian matrix, i.e. such that $L_{aa}=\sum_{b\neq a}|L_{ab}|\geq 0$ and $L_{ab}=L_{ba}\leq 0$, for all $a\neq b$. The matrix $L$ can be thought as the Laplacian of a graph encoding (prior) similarities between alternatives, the weight $|L_{ab}|$ representing the similarity between $a$ and $b$. Therefore, the regularization term $\sum_{ab}\theta_{a}L_{ab}\theta_{b}=\frac{1}{2}\sum_{a\neq b}|L_{ab}|(\theta_{a}-\theta_{b})^{2}$ incentivizes the model to (a priori) assign similar scores to similar alternatives.
We can now define the class of GBT models that we will study in this paper.
Definition 3 (Linear GBT with Diffusion Prior).
Let $f$ be a root law, $x$ be an embedding, $\sigma>0$ a positive constant, and $L$ a Laplacian matrix. The model $\textsc{GBT}_{f,\sigma,x,L}$ is defined as
$\textsc{GBT}_{f,\sigma,x,L}({\bf D})\triangleq x^{T}\beta^{*}({\bf D})\in\mathbb{R}^{A},$
where $\beta^{*}({\bf D})\triangleq\operatorname*{arg\,min}\mathcal{L}(\cdot|{\bf D})$ minimizes the strongly convex loss function
$\mathcal{L}(\beta|{\bf D})=\mathcal{R}(\beta)+\sum_{(a,b,r)\in{\bf D}}\Phi_{f}(x_{ab}^{T}\beta)-rx_{ab}^{T}\beta.$ (4)
Remark that the original GBT is a special case of Linear GBT with Diffusion Prior with $A=D$, $x=I$ the identity matrix, and $L=0$.
Proposition 2.
Linear GBT with diffusion prior is neutral, i.e. invariant up to alternative relabeling.
Proof.
See Appendix A for a formal statement and derivation. ∎
3.2 Monotonicity and diffusion
We now present our main result (Theorem 1). We prove that for a special class of embeddings, namely diffusion embeddings, monotonicity is guaranteed. Diffusion embeddings take their name from the interplay with (super) laplacian matrices.
Definition 4 (Super-Laplacian matrix).
A super-Laplacian matrix $\Delta$ is a symmetric matrix such that for all $a\neq b$, $\Delta_{aa}>-\sum_{b\neq a}\Delta_{ab}$ and $\Delta_{ab}\leq 0$.
Definition 5 (Diffusion embedding).
An embedding $x$ is a diffusion embedding if the Gram matrices $X_{\lambda}=x^{T}x+\lambda I$ have super-Laplacian inverses $X_{\lambda}^{-1}$ for any $\lambda>0$.
Note that if $X=x^{T}x$ is itself invertible with super-Laplacian inverse, then it is a diffusion embedding. However, this case is restrictive since it implies that $D\geq A$.
Theorem 1 (Monotonicity with diffusion embeddings).
For any root law $f$, positive constant $\sigma>0$, diffusion embedding $x$, and Laplacian matrix $L$, $\textsc{GBT}_{f,\sigma,x,L}$ is monotone.
Proof.
The theorem follows directly from Proposition 3 and Theorem 3, which are provided and proved in Section 4. ∎
3.3 Example: one-hot encoding
A one-hot encoding is possible when the alternatives can be arranged into multiple disjoint classes. For example, if the alternatives represent videos on YouTube, one can partition them by the YouTube channel they belong to. In that case, the score of an alternative $a$ is defined as $\theta_{a}=\gamma_{d(a)}+s^{2}\cdot\alpha_{a}$, where $d(a)$ is the channel of $a$. The score $\gamma_{d(a)}$ represents the score of the channel $d(a)$, while $\alpha_{a}$ represents a residual score of $a$, and $s$ is a real constant that controls the scale of the residual score. Theorem 2 states a one-hot encoding is an example of diffusion embedding. We postpone the proof to Appendix G.
Theorem 2 (GBT with one-hot encoding).
Let $f$ be a root law, $\sigma>0$ a positive constant, $L$ a Laplacian matrix and $s\in\mathbb{R}$. Let $x:\mathbb{R}^{D\times\mathcal{A}}$ be a one-hot encoding matrix: $x_{da}=1$ iff $a$ belongs to $d$. Then, for any real number $s$, $\begin{pmatrix}x&sI\end{pmatrix}^{T}$ is a diffusion embedding and the score $\textsc{GBT}_{f,\sigma,x,L}$ is monotone.
4 The proof
This Section provides the mathematical analysis that builds to the proof of the main result, Theorem 1. Section 4.1 propose a differential analysis framework for the dataset operations outlined above; Section 4.2 then provide the proof.
4.1 Differential analysis of dataset operations
[…]
Because the loss function is a sum of terms indexed by the elements of the dataset, this relation is quite simple
$\mathcal{L}(\beta|\operatorname*{\textsc{append}}_{a,b,r}({\bf D}))$ $=\mathcal{L}(\beta|{\bf D})+\Phi_{f}(\theta_{ab}(\beta))-r\theta_{ab}(\beta)$ (5)
$\mathcal{L}(\beta|\operatorname*{\textsc{update}}_{n,r}({\bf D}))$ $=\mathcal{L}(\beta|{\bf D})-(r-r_{n})\theta_{a_{n}b_{n}}(\beta)$ (6)
To enable the differential analysis of these operations, we introduce a smooth deformation of the loss function.
Definition 6 (Smoothened loss).
For every $\lambda\in\mathbb{R}$, and every operation $o$ of the form $\operatorname*{\textsc{append}}(c)$ or $\operatorname*{\textsc{update}}(n,c)$, we define the smoothened loss $\mathcal{L}_{\lambda}$ by
$\mathcal{L}_{\lambda}(\beta|{\bf D},o)\triangleq\mathcal{L}(\beta|{\bf D})+\lambda\cdot\begin{cases}\Phi_{f}(\theta_{ab}(\beta))-r\theta_{ab}(\beta)&\text{if }o=\operatorname*{\textsc{append}}(a,b,r),\\ -(r-r_{n})\cdot\theta_{a_{n}b_{n}}(\beta)&\text{if }o=\operatorname*{\textsc{update}}(n,r),\\ 0&\text{otherwise.}\end{cases}$ (7)
Denote also $\beta^{*}_{\lambda}({\bf D},o)\triangleq\operatorname*{arg\,min}\mathcal{L}_{\lambda}(\cdot|{\bf D},o)$ and $\theta^{*}_{\lambda}({\bf D},o)\triangleq\theta(\beta^{*}_{\lambda}({\bf D},o))$.
The smoothened loss matches the loss at $\lambda\in\left\{0,1\right\}$, as $\mathcal{L}_{0}(\beta|{\bf D},o)=\mathcal{L}(\beta|{\bf D})$ and $\mathcal{L}_{1}(\beta|{\bf D},o)=\mathcal{L}(\beta|o({\bf D}))$. We will leverage this by using the integral expression
[…]
When $\lambda\mapsto\theta^{*}_{\lambda}({\bf D},o)$ is continuously differentiable, this integral expression is well-defined, and it suffices that the derivative $d\theta^{*}_{\lambda a}({\bf D},o)/d\lambda$ be non-negative for the score difference at $a$ to be non-negative. Lemma 1 states that this derivative is well defined and provides a formula. The proof is given in Appendix B
Lemma 1.
Let $H=H(\theta|{\bf D})$ denote the Hessian of $\mathcal{E}(\theta|{\bf D})=\sum_{(a,b,r)\in{\bf D}}\Phi_{f}(\theta_{ab})$. We write $X=\sigma^{2}x^{T}x$ for the (scaled) Gram matrix of the embedding $x$. Then, for any basic operation $o$ and dataset ${\bf D}$, the loss function $\mathcal{L}(\beta|o_{\lambda}({\bf D}))$ admits a unique global minimizer $\beta^{*}_{\lambda}({\bf D},o)$, and the inferred score $\theta^{*}_{\lambda}({\bf D},o)$ is a smooth function of $\lambda$ over $[0,\infty)$. Moreover, if $o=\operatorname*{\textsc{update}}(n,r)$, ${\bf D}_{n}\simeq(a,b,s)$, then the score of the alternative $a$ satisfies
$\frac{d\theta^{*}_{\lambda a}({\bf D},o)}{d\lambda}\bigg{|}_{\lambda=\mu}=(r-s)\cdot e_{a}^{T}\big{(}I+X(L+H)\big{)}^{-1}Xe_{ab}.$ (9)
Similarly, if $o=\operatorname*{\textsc{append}}(a,b,r)$
$\frac{d\theta^{*}_{\lambda a}({\bf D},o)}{d\lambda}\bigg{|}_{\lambda=\mu}=(r-\Phi^{\prime}_{f}(\theta^{*}_{ab}))\cdot e_{a}^{T}\big{(}I+X(L+H+\mu\Phi_{f}^{\prime\prime}(\theta_{ab})\cdot S_{ab})\big{)}^{-1}Xe_{ab},$ (10)
where $S_{ab}$ is the Laplacian of the graph over the alternatives with a single edge $ab$ (with weight $1$), i.e. $S_{ab}[a,a]=S_{ab}[b,b]=1$, $S_{ab}[a,b]=S_{ab}[b,a]=-1$, and $S_{ab}[c,d]=0$ otherwise.
We are interested in the sign (positive or negative) of the expressions in Equations (9) and (10). First, the factors $r-s$ and $r-\Phi^{\prime}_{f}(\theta^{*}_{ab})$ are easy to understand. If $o=\operatorname*{\textsc{update}}(n,r)$ favors $a$ over $b$ then $r-s\geq 0$ by definition. If $o=\operatorname*{\textsc{append}}(a,b,r)$ favors $a$ over $b$, then $r=\sup\mathcal{R}$ which, by (DBLP:conf/aaai/FageotFHV24, , Theorem 1), is known to be the supremum of $\Phi^{\prime}_{f}$.
Therefore, if we want to compare the scores of $a$ and $b$, the important factor to study is the matrix $(I+X(L+\tilde{H}))^{-1}X$, where $\tilde{H}=H+\mu\Phi_{f}^{\prime\prime}(\theta_{ab})\cdot S_{ab}$ if $o=\operatorname*{\textsc{append}}(a,b,r)$, or $\tilde{H}=H$ if $o=\operatorname*{\textsc{update}}(n,(a,b,r))$. We study this matrix in the next section.
4.2 Good embeddings: a sufficient condition for monotonicity
In this Section, we provide a sufficient condition on the embedding for the Linear GBT model to be monotone (Definition 7, Theorem 3), and provide mathematical properties of this condition (Proposition 9). Finally, we show that diffusion embeddings meet the above sufficient condition (Proposition 3).
Definition 7 (Good embeddings).
Given a Laplacian matrix $Y$, an embedding $x$ is $Y$-good if the Gram matrix $X=x^{T}x$ satisfies $e_{a}^{T}(I+XY)^{-1}Xe_{ab}\geq 0$ for all $(ab)$. An embedding $x$ is good if $x$ is $Y$-good for all Laplacian matrices $Y$.
Theorem 3 (Monotonicity with good embeddings).
For any root law $f$, positive constant $\sigma>0$, Laplacian matrix $L$, and good embedding $x$, $\textsc{GBT}_{f,\sigma,x,L}$ is monotone.
Before going to the proof, we provide some intuition. Note first that, given a Laplacian matrix $L$, the Hessian $H$ of $\mathcal{E}(\theta|{\bf D})=\sum_{(a,b,r)\in{\bf D}}\Phi_{f}(\theta_{ab})$ is also a Laplacian matrix. Indeed, let $G$ be the weighted graph whose edges are the pairs $(ab)$ of alternatives that occur in the dataset ${\bf D}$, weighted by $G_{ab}=N_{ab}\cdot\Phi^{\prime\prime}_{f}(\theta_{ab})$ where $N_{ab}$ is the number of occurrences of the pair $(ab)$ in ${\bf D}$. Since $\Phi_{f}$ is convex, these weights are nonnegative. Then, a direct calculation shows that $H$ is the graph Laplacian of the weighted graph $G$: for $a\neq b$, $H_{aa}=\sum_{c\neq a}N_{ac}\Phi_{f}^{\prime\prime}(\theta_{ac})$ and $H_{ab}=-N_{ab}\Phi_{f}^{\prime\prime}(\theta_{ab})$. Therefore, the matrix $L+\tilde{H}$, where $\tilde{H}$ is defined in section 4.1, is the Laplacian of a graph that combines the prior similarities ($L$) with the similarities inferred from the dataset at hand ($\tilde{H}$). This observation motivates Definition 7.
Proof of Theorem 3.
By Lemma 1, if, for every operation $o$, every score function $\theta$ and every dataset ${\bf D}$, the inequality $e_{a}^{T}(I+X(L+\tilde{H}))^{-1}Xe_{ab}\geq 0$ holds, then the score $\theta^{*}({\bf D})$ derived from the loss function of Equation (4) is monotone. This is precisely implied by $x$ being $L$-good. ∎
This result motivates a more precise understanding of good embeddings. In the special cases $(A,D)=(2,D)$ and $(A,D)=(A,1)$, we have complete characterizations (see Appendix C). In general, however, checking goodness is not straightforward: two embeddings $x$ and $y$ can be individually good, while their concatenation $\begin{bmatrix}x&y\end{bmatrix}^{\top}$ fails to be good (see Proposition 8, Appendix D). Nonetheless, any embedding can be made $Y$-good by concatenating it with a sufficiently scaled identity. We formalize this in Appendix E.
4.3 Diffusion embeddings are good
Finally, we show that any diffusion embedding is a good embedding. This uses the fact that any super-Laplacian matrix $\Delta$ satisfies $e_{a}^{T}\Delta^{-1}e_{ab}\geq 0$ for any pair $ab$. This result has been proved in (DBLP:conf/aaai/FageotFHV24, , Lemma 1) and we provide an alternative proof highlighting the diffusion perspective F.
Proposition 3.
Any diffusion embedding is a good embedding.
Proof.
If $x$ is a diffusion embedding, $\lambda>0$, and $Y$ is an arbitrary Laplacian matrix, then the matrix $X_{\lambda}^{-1}+Y$, where $X_{\lambda}=x^{T}x+\lambda I$, is super-Laplacian. Consequently, $e_{a}^{T}(I+X_{\lambda}Y)^{-1}X_{\lambda}e_{b}=e_{a}^{T}(X_{\lambda}^{-1}+Y)^{-1}e_{b}\geq 0$. The claim follows by taking the limit $\lambda\to 0$. ∎
5 Experimental evaluation
In this Section, we provide a numerical exploration of the prevalence of “goodness” for random embeddings, and the statistical error of several preference learning models. [^3] The code is available as supplementary material, and will be made publicly after the review process. We run experiments on a personal laptop with 16GB of RAM and a 2.10 GHz processor.
5.1 Probability of goodness for random embeddings
To illustrate the challenges of achieving good embeddings, we generate random i.i.d. Gaussian embedding matrices x and evaluate their quality. In Figure 1, we examine a single Gaussian embedding $x$ (left) and its concatenation with the identity matrix $I$ (right). Our findings indicate that the goodness of $x$ is more likely for large values of $D/A$ and significantly diminishes when $A/D$ is large. The concatenation with the identity matrix notably enhances the goodness, aligning with Proposition 9.
Figure 1: Left pane: Probability that a Gaussian i.i.d embedding $x$ is a good embedding for $2\leq A\leq 15$ and $1\leq D\leq 15$. Right pane: As for the left pane with embedding $\begin{bmatrix}I&x\end{bmatrix}^{T}$.
5.2 Generative model, metric, and simulations
Figure 2: Left pane: nMSE as a function of $D$ for $A=25$ alternatives and $N=500$ comparisons over $100$ seeds. Blue curve with $\begin{bmatrix}I&x\end{bmatrix}^{T}$ (full embedding), orange curve with embedding $I$ (classical GBT), and green curve with embedding $x$ (features only). Right pane: nMSE with respect to the number of comparisons $N$ for $A=20$, $D=10$, and $1000$ seeds. Blue curve: GBT with one-hot encoding; Orange curve: GBT.
For each experiment, we consider the ground-truth embedding $x^{\dagger}\in\mathbb{R}^{D\times A}$, Laplacian matrix $L^{\dagger}$, constant $\sigma^{\dagger}$, and root law $f^{\dagger}$. The ground-truth features are generated as $\beta^{\dagger}\sim\mathcal{N}(0,(\sigma^{\dagger})^{2}I+x^{\dagger}L^{\dagger}(x^{\dagger})^{T})$ and $\theta^{\dagger}=(x^{\dagger})^{T}\beta^{\dagger}$. We then create a dataset ${\bf D}:[N]\rightarrow[A]^{2}\times\mathcal{R}$ by first selecting $N$ random comparison pairs uniformly. The corresponding random comparisons $r$ are generated using the root law $f^{\dagger}$ and conditionally to $\theta^{\dagger}$. We shall only consider the uniform root law $f^{\dagger}=\frac{1}{2}1_{[-1,1]}$ and set $\sigma^{\dagger}=1$.
The estimated scores are computed as $\theta^{*}({\bf D})=\text{GBT}_{f,\sigma,x,L}({\bf D})$, where the quadruplet $(f,\sigma,x,L)$ may or may not align with the ground truth. Since the quality of a score vector is invariant under constant shifts, we evaluate the error over zero-mean versions of both $\theta^{\dagger}$ and $\theta^{*}({\bf D})$. More precisely, we use Monte Carlo simulations to estimate the normalized mean squared error (nMSE), defined as:
$\text{nMSE}(f^{\dagger},\sigma^{\dagger},x^{\dagger},L^{\dagger};f,\sigma,x,L;N)=\text{nMSE}=\mathbb{E}\left[\frac{\left\|\left(\theta^{*}({\bf D})-\bar{\theta}^{*}({\bf D})\right)-\left(\theta^{\dagger}-\bar{\theta}^{\dagger}\right)\right\|^{2}}{\|\theta^{\dagger}-\bar{\theta}^{\dagger}\|^{2}}\right].$
We then analyze how the nMSE evolves with respect to various parameters.
Figure 2(a) shows the nMSE as a function of $D$, using data generated with $x^{\dagger}=\begin{bmatrix}I&\tilde{x}^{\dagger}\end{bmatrix}^{T}$ (i.i.d. Gaussian $\tilde{x}$), uniform $f^{\dagger}$, $L^{\dagger}=0$, and $\sigma^{\dagger}=1$. We compare three models with shared parameters $(f,L,\sigma)=(f^{\dagger},L^{\dagger},\sigma^{\dagger})$, using $x^{\dagger}$, $I$ (classical GBT), and $\tilde{x}^{\dagger}$ respectively. The embedding-based model outperforms others, combining the strengths of classical GBT for small $D$ and feature-based learning for larger $D$.
Figure 2(b) shows the nMSE as a function of the number of comparisons $N$. Data are generated with $(f^{\dagger},x^{\dagger},L^{\dagger},\sigma^{\dagger})=\left(\frac{1}{2}1_{[-1,1]},\begin{bmatrix}I&\tilde{x}^{\dagger}\end{bmatrix}^{T},0,1\right)$, where $\tilde{x}$ is a one-hot encoding matrix (see Section 3.3). The results show that one-hot encoding greatly reduces the number of comparisons needed to reach a given nMSE. This is useful in applications like YouTube score estimation hoang2024solidagomodularcollaborativescoring , where the encoding reflects the channel and enables generalization across alternatives.
6 Conclusion
In this paper, we introduced a new comparison-based preference learning model, namely linear GBT with diffusion prior. This model not only generalizes to previously uncompared data using embeddings, but also potentially guarantees monotonicity, depending on the class of embeddings used. We proved that our model is monotone for various classes of embeddings (one-hot encodings, diffusion, and good embeddings). To the best of our knowledge, linear GBT with diffusion prior is the first model that guarantees monotonicity while being able to generalize.
Diffusion embeddings form a class of embeddings (containing one-hot encodings) that yield monotonicity. Our proof techniques relied on an interesting interplay between an algebraic criterion for monotonicity (Definition 7) and properties of (super) Laplacian matrices akin to diffusion theory.
Limitations.
While improving the understanding of preference learning with guarantees, our theory currently provides guarantees for diffusion embeddings only. We hope to motivate more work on preference learning with guarantees, in order to build more trusworthy AI systems, with notable applications in weyl2025prosocialmedia ; hoang2021tournesolquestlargesecure ; hoang2024solidagomodularcollaborativescoring ; small2021polis . Also, we caution against the use of preference learning algorithms that rely on data collected in inhumane conditions, as is mostly the case today DW2025 ; perrigo2023exclusive ; hao2023cleaning ; ghanalawsuit . It is unclear whether our work can positively contribute to this issue.
Appendix A Proof of neutrality (Proposition 2)
To formalize neutrality, we must define how a permutation $\tau$ of the alternatives acts on the inputs and outputs of Linear GBT with Diffusion Prior. We define the actions as follow:
$(\tau\cdot{\bf D})_{n}$ $=\tau\cdot{\bf D}_{n},$ (11)
$\tau\cdot(a,b,r)$ $=(\tau(a),\tau(b),r),$ (12)
$(\tau\cdot\theta)_{a}$ $=\theta_{\tau(a)},$ (13)
$(\tau\cdot x)_{a}$ $=x_{\tau(a)},$ (14)
$(\tau\cdot L)_{ab}$ $=L_{\tau(a)\tau(b)}.$ (15)
Neutrality is formalized by the equality
$\forall\tau,~{}\tau^{-1}\cdot GBT_{f,\sigma,\tau\cdot x,\tau\cdot L}\circ\tau=GBT_{f,\sigma,x,L}.$ (16)
We indeed have
$\left(\tau^{-1}\cdot GBT_{f,\sigma,\tau\cdot x,\tau\cdot L}\circ\tau({\bf D})\right)_{a}=\left(GBT_{f,\sigma,\tau\cdot x,\tau\cdot L}\left(\tau\cdot{\bf D}\right)\right)_{\tau^{-1}(a)}$ (17)
$=(\tau\cdot x)^{T}_{\tau^{-1}(a)}\beta_{f,\sigma,\tau\cdot x,\tau\cdot L}^{*}(\tau\cdot{\bf D})=x_{a}^{T}\beta_{f,\sigma,\tau\cdot x,\tau\cdot L}^{*}(\tau\cdot{\bf D})$ (18)
$=x^{T}_{a}\operatorname*{arg\,min}_{\beta}\frac{1}{2\sigma^{2}}\left\lVert{\beta}\right\rVert_{2}^{2}+\frac{1}{2}\sum_{ab}(x^{T}_{\tau(a)}\beta)L_{\tau(a)\tau(b)}(x^{T}_{\tau(b)}\beta)$
$\qquad+\sum_{(a,b,r)\in{\bf D}}\left(\Phi_{f}((\tau\cdot x)_{\tau^{-1}(a)\tau^{-1}(b)}^{T}\beta)-r(\tau\cdot x)_{\tau^{-1}(a)\tau^{-1}(b)}^{T}\beta\right)$ (19)
$=x^{T}_{a}\operatorname*{arg\,min}_{\beta}\frac{1}{2\sigma^{2}}\left\lVert{\beta}\right\rVert_{2}^{2}+\frac{1}{2}\sum_{a^{\prime}b^{\prime}}(x^{T}_{a^{\prime}}\beta)L_{a^{\prime}b^{\prime}}(x^{T}_{b^{\prime}}\beta)+\sum_{(a,b,r)\in{\bf D}}\left(\Phi_{f}(x_{ab}^{T}\beta)-rx_{ab}^{T}\beta\right)$ (20)
$=x_{a}^{T}\beta_{f,\sigma,x,L}^{*}({\bf D})=\left(GBT_{f,\sigma,x,L}({\bf D})\right)_{a}.$ (21)
Appendix B Proof of Lemma 1
We consider the case of $o=\operatorname*{\textsc{append}}(a,b,r)$. The other case, $o=\operatorname*{\textsc{update}}(n,(a,b,r))$, is proved similarly. The loss functions $\mathcal{L}_{\lambda}(\beta|{\bf D},o)$ and $\mathcal{L}(\beta|{\bf D})$ differ by $\Phi_{f}(\theta_{ab})-r\theta_{ab}$. The term $\theta_{ab}$ being linear in $\beta$, its Hessian is zero. On the other hand, the Hessian of $\Phi_{f}(\theta_{ab})$ is simply the Laplacian $\Phi_{f}^{\prime\prime}(\theta_{ab})\cdot S_{ab}$ of the graph with a single edge $ab$ with weight $\Phi_{f}^{\prime\prime}(\theta_{ab})$. Writing
$H^{\lambda}=H+\lambda\Phi_{f}^{\prime\prime}(\theta_{ab})\cdot S_{ab},$ (22)
we obtain the Hessian of the loss
$H_{\beta}\mathcal{L}_{\lambda}(\beta|{\bf D},o)=x(L+H^{\lambda})x^{T}+\frac{1}{\sigma^{2}}I\succeq\frac{1}{\sigma^{2}}I.$ (23)
Therefore, the loss $\mathcal{L}_{\lambda}(\beta|{\bf D},o)$ is strictly convex, and admits a global minimizer $\beta^{*}_{\lambda}({\bf D},o)$. To simplify notations in this proof, we write $\beta^{*}(\lambda)$.
We want to use the implicit function theorem to analyze how $\beta^{*}(\lambda)$ varies with $\lambda$. For that, consider the gradient of the loss $\mathcal{L}_{\lambda}(\beta|{\bf D},o)$
$F(\beta,\lambda)=\nabla_{\beta}\mathcal{L}(\beta|o_{\lambda}({\bf D}))=\nabla_{\beta}\mathcal{L}(\beta|{\bf D})+\lambda\cdot(\Phi^{\prime}_{f}(\theta_{ab})-r)xe_{ab}$ (24)
with domain $\mathbb{R}^{D}\times\mathbb{R}$ and codomain $\mathbb{R}^{D}$.
Fix some $\mu\in\mathbb{R}$. The Jacobian of $F(\beta,\lambda)$ with respect to $\beta$, evaluated at $\lambda=\mu$, is given by
$(J_{\beta}F)_{|\lambda=\mu}=\frac{1}{\sigma^{2}}I+x(L+H^{\mu})x^{T}$ (25)
which is invertible. Moreover, since $\beta^{*}(\mu)$ is a minimizer, we have $F(\beta^{*}(\mu),\mu)=0$.
Hence, the implicit function theorem states that there exists an open neighborhood $U$ of $\mu$, and a smooth function $\gamma:U\to\mathbb{R}^{D}$ such that
$\gamma(\mu)=\beta^{*}(\mu)$ (26)
$\forall\lambda\in U,F(\gamma(\lambda),\lambda)=0$ (27)
The latter equality implies that $\gamma(\lambda)=\beta^{*}(\lambda)$ for all $\lambda\in U$. Consequently, $\beta^{*}(\lambda)$ and $\theta^{*}(\lambda)=x^{T}\beta^{*}(\lambda)$ depend smoothly on $\lambda$. In addition, the implicit function theorem also gives an expression for the Jacobian $J_{\lambda}\beta^{*}$, evaluated at $\mu$
$(J_{\lambda}\beta^{*})_{|\lambda=\mu}$ $=-(J_{\beta}F)^{-1}J_{\lambda}F$ (28)
$=(r-\Phi^{\prime}_{f}(\theta^{*}_{ab}))\cdot\big{(}\frac{1}{\sigma^{2}}I+x(L+H^{\mu})x^{T}\big{)}^{-1}xe_{ab}$ (29)
Finally, note that
$\frac{d\theta^{*}(\lambda)}{d\lambda}_{|\lambda=\mu}$ $=J_{\beta}\theta\cdot J_{\lambda}\beta^{*}$ (30)
$=(r-\Phi^{\prime}_{f}(\theta^{*}_{ab}))\cdot x^{T}\big{(}\frac{1}{\sigma^{2}}I+x(L+H^{\mu})x^{T}\big{)}^{-1}xe_{ab}$ (31)
Let $M=L+H^{\mu}$, and $X=\sigma^{2}x^{T}x$. Using Woodbury’s identity $(I+UV)^{-1}=I-U(I+VU)^{-1}V$, we derive
$\big{(}\frac{1}{\sigma^{2}}I+x(L+H^{\mu})x^{T}\big{)}^{-1}$ $=\sigma^{2}(I+\sigma^{2}xMx^{T})^{-1}$ (32)
$=\sigma^{2}I-\sigma^{4}xM(I+\sigma^{2}x^{T}xM)^{-1}x^{T}$ (33)
$x^{T}\big{(}\frac{1}{\sigma^{2}}I+x(L+H)x^{T}\big{)}^{-1}x$ $=X-XM(I+XM)^{-1}X$ (34)
$=(I-XM(I+XM)^{-1})X$ (35)
$=(I+XM)^{-1}X$ (36)
The result follows.
Appendix C Good Embeddings for $A=2$ or $D=1$
We can fully characterize goodness in lowest dimensional regime, either with only two alternatives or with an embedding on a single feature.
Definition 8.
We say that a matrix $M$ is max-diagonally dominant if $M_{aa}\geq M_{ab}$ for any $(ab)$.
Proposition 4.
An embedding with Gram matrix $X=\begin{bmatrix}a&c\\ c&b\end{bmatrix}$ is monotonicity proof if and only if $-\sqrt{ab}\leq c\leq\min(a,b)$.
Proof.
We first observe that, since $x^{T}x$ is positive semidefinite, $a,b\geq 0$ and $det(x^{T}x)=ab-c^{2}\geq 0$. This implies in particular that $c\geq-\sqrt{ab}$, which is the desired lower bound.
Two-dimensional Laplacian matrices are of the form $Y=\delta\left[\begin{matrix}1&-1\\ -1&1\end{matrix}\right]$ with $\delta>0$, which can be chosen as being equal to $1$ without loss of generality. We then have that $I+XY=\left[\begin{matrix}1+a-c&c-a\\ c-b&1+b-c\end{matrix}\right]$. After simplification, its determinant is $det(I+XY)=1+a+b-2c$. This quantity is strictly positive for $c<\frac{1+a+b}{2}$, which is always the case for as $c^{2}\leq ab$. Hence, the matrix is invertible and we have, after computation,
$M=(I+XY)^{-1}X=\frac{1}{1+a+b-2c}\left[\begin{matrix}a+ab-c^{2}&c+ab-c^{2}\\ c+ab-c^{2}&b+ab-c^{2}.\end{matrix}\right]$ (37)
Then, $M$ is max-diagonally dominant if and only if $a\geq c$ et $b\geq c$, as expected. Finally, this shows that the embedding $x$ is $Y$-good for any $Y$, hence good. ∎
We now focus on the case $A=2$, for which we obtain a complete characterization of the goodness.
Proposition 5.
Consider the GBT model with embedding $x=[x_{a},x_{b}]\in\mathbb{R}^{D\times 2}$ and root law $f$. Then, the model is good if and only if, for any $(ab)$, $x_{a}=x_{b}$ or $x_{a}^{T}x_{b}\leq min(\|x_{a}\|^{2},||x_{b}\|^{2})$. The latter is equivalent to
$\theta(x_{a},x_{b})\leq\theta_{0}(\|x_{a}\|,\|x_{b}\|)=\arccos\left(\min\left(\frac{\|x_{a}\|}{\|x_{b}\|},\frac{\|x_{b}\|}{\|x_{a}\|}\right)\right)\in[0,\pi/2]$ (38)
where $\theta(x_{a},x_{b})=\frac{x_{a}^{T}x_{b}}{\|x_{a}\|\|x_{b}\|}$ is the angle between $x_{a}$ and $x_{b}$.
Proof.
We apply Proposition 4 to $a=x_{a}^{2}$, $b=x_{b}^{2}$, and $c=x_{a}x_{b}$. Then, the goodness is equivalent to $x_{a}x_{b}\leq\min(x_{a}^{2},x_{b}^{2})$. This relation is true for $x_{a}=x_{b}$ and $x_{a}x_{b}<0$. Otherwise, we have $0<x_{a},x_{b}$. The relations $x_{a}x_{b}\leq x_{a}^{2}$ and $x_{a}x_{b}\leq x_{b}^{2}$ implies that $x_{b}\leq x_{a}$ and $x_{a}\leq x_{b}$ respectively, leading to a contradiction. Finally, the goodness is equivalent to $x_{a}=x_{b}$ or $x_{a}$ and $x_{b}$ have different signs.
For the $D$ dimensional case, the same observation holds with $a=\|x_{a}\|^{2}$, $b=\|x_{b}\|^{2}$, and $c=x_{a}^{T}x_{b}$. Then, Proposition 4 implies that the goodness is equ bivalent to $x_{a}^{T}x_{b}\leq\min(\|x_{a}\|^{2},\|x_{b}\|^{2})$. The angular characterization follows easily. ∎
We shall see that the goodness is very restricted for $D=1$.
Proposition 6.
We consider the GBT model with embedding $x=[x_{a}]_{a\in\mathcal{A}}\in\mathbb{R}^{1\times A}$, root law $f$, and Laplacian matrix $L=0$. The model is good if and only if
$x=[u,\ldots,u,-v,\ldots,-v,0,\ldots,0]^{T}P=[u1_{A_{1}},-v1_{A_{2}},0_{A_{3}}]^{T}P$ (39)
for some $u,v>0$ and $P$ a permutation matrix.
Proof.
The goodness is equivalent to the fact that, for any $(ab)$, $(x_{a}-x_{b})x_{a}\geq 0$ and $(x_{b}-x_{a})x_{b}\geq 0$. This is equivalent to $x_{a}x_{b}\leq min(x_{a}^{2},x_{b}^{2})$, i.e. $x_{a}=x_{b}$ or $x_{a}x_{b}\leq 0$. This means that any $x_{a}>0$ should have a common value $u>0$ and any $x_{b}<0$ should have a common value $-v<0$. Hence, $x_{a}$ can take only the values $u,-v$ and 00. Permuting the values, we obtain (39). ∎
Appendix D Counterexamples for Monotonicity
Proposition 7.
There exists good embeddings $x_{1}$ and $x_{2}$ such that $x=\begin{bmatrix}x_{1}&x_{2}\end{bmatrix}^{T}$ is not good.
Proof.
For $A=3$ and a Gaussian root law ($Y=3I-J$), we consider $x_{1}$ and $x_{2}$ such that
$X_{1}=\begin{bmatrix}0&0&0\\ 0&1&1\\ 0&1&1\end{bmatrix}\quad\text{and}\quad X_{2}=\begin{bmatrix}1&1&0\\ 1&1&0\\ 0&0&0\end{bmatrix}$
Then, $X=x^{T}x=x_{1}^{T}x_{1}+x_{2}^{T}x_{2}=X_{1}+X_{2}=\begin{bmatrix}1&1&0\\ 1&2&1\\ 0&1&1\end{bmatrix}$. Then, $x_{1}$ and $x_{2}$ are good embedding (e.g. remarking that they are $J$-blocs matrices and using Theorem 2). The matrix $M=(I+XY)^{-1}X$ is given by
$M=\frac{1}{8}\begin{bmatrix}{3}&{4}&1\\ {4}&8&4\\ 1&4&3\end{bmatrix}$
and verifies $M_{12}>M_{11}$. This contradicts Definition 7 and $x$ is not $Y$-proof for $Y=3I-J$, therefore not good. ∎
Proposition 8.
There exists one-hot encodings $x_{1}$ and $x_{2}$ such that $x=\begin{bmatrix}x_{1}&x_{2}\end{bmatrix}^{T}$ is not a good embedding.
Proof.
For $A=5$, let $x_{1}$ and $x_{2}$ be one-hot encoding with Gram matrices
$X_{1}=\begin{bmatrix}1&1&1&0&0\\ 1&1&1&0&0\\ 1&1&1&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\end{bmatrix}\quad\text{and}\quad X_{2}=\begin{bmatrix}1&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&1&1\\ 0&0&1&1&1\\ 0&0&1&1&1\end{bmatrix}.$
Then, $x_{1}$ and $x_{2}$ are good embeddings according to Theorem 2. The Gram matrix of the concatenated embedding $x=\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}$ is
$X=X_{1}+X_{2}=\begin{bmatrix}2&1&1&0&0\\ 1&2&1&0&0\\ 1&1&2&1&1\\ 0&0&1&2&1\\ 0&0&1&1&2\end{bmatrix}$
is not a $Y$-good embedding for the Laplacian matrix
$Y=\begin{pmatrix}2&-1&0&0&-1\\ -1&4&-1&-1&-1\\ 0&-1&1&0&0\\ 0&-1&0&1&0\\ -1&-1&0&0&2\end{pmatrix}.$
We indeed have that
$M=(I+XY)^{-1}X\approx\begin{pmatrix}0.96&0.70&0.85&0.52&0.67\\ 0.70&0.90&0.95&0.67&0.68\\ 0.85&0.95&1.48&0.83&0.84\\ 0.52&0.67&0.83&1.08&0.56\\ 0.67&0.68&0.84&0.56&0.93\end{pmatrix}$
is such that $M_{23}>M_{22}$. ∎
Appendix E Results of Section 4.2
We now show that any embedding can be made $Y$-good by appending a sufficiently dominant identity component. This result guarantees that, asymptotically, adding uncorrelated features improves embedding monotonicity, regardless of the original embedding structure. This can be made in relation with Figure 1 for which we compare i.i.d. Gaussian $x$ with it’s concatenation with $I$.
Proposition 9.
For any embedding $x$ and any Laplacian matrix $Y$, the embedding $x_{\lambda}=\begin{bmatrix}I&x/\lambda\end{bmatrix}^{T}$ is $Y$-good for any $\lambda>3\sqrt{A}\|x^{T}x\|/\mathrm{DiagDom}(\sigma^{2}Y)$ where $\mathrm{DiagDom}(Y)=\min_{(ab)}(I+Y)^{-1}_{aa}-(I+Y)^{-1}_{ab}>0$.
Proof of Proposition 9.
We set $M(X,Y)=(I+XY)^{-1}X$. The Frobenius norm is such that $\|X+Y\|\leq\|X\|+\|Y\|$ and $\|XY\|\leq\|X\|\|Y\|$. Assuming that $\|X\|<1$ and for $Z=(I+X)^{-1}-(I-X)$, we have
$\|Z\|\leq\|X\|^{2}\sum_{k\geq 0}\|X\|^{k}=\frac{\|X\|^{2}}{1-\|X\|}.$ (40)
The matrix $I+X/\lambda$ is positive definite, hence invertible and we have
$M(I+X/\lambda,Y)=(I+(I+X/\lambda)Y)^{-1}(I+X/\lambda)=((I+X/\lambda)^{-1}+Y)^{-1}.$ (41)
Then, there exist matrices $Z$ and $W$, that we will control later on, such that
$M(I+X/\lambda,Y)$ $=(I-X/\lambda+Z+Y)^{-1}$ (42)
$=(I+Y)^{-1}(I+(Z-X/\lambda)(I+Y)^{-1})^{-1}$ (43)
$=(I+Y)^{-1}(I+(Z-X/\lambda)(I+Y)^{-1}+W).$ (44)
From now, we assume that $\lambda>3\sqrt{A}\|X\|$. In particular, using that $h:x\mapsto\frac{x}{1-x}$ is increasing over $(0,1)$ we have that $\frac{\|X\|/\lambda}{1-\|X\|/\lambda}\leq h(1/3\sqrt{A})\leq h(1/3)=\frac{1}{2}$. According to (45) applied to $X/\lambda$, we therefore have have
$\|Z\|\leq\frac{\|X/\lambda\|^{2}}{1-\|X/\lambda\|}\leq\frac{\|X\|}{2\lambda}.$ (45)
We also have that
$\|X/\lambda-Z\|\leq\frac{\|X\|}{\lambda}+\frac{\|X\|}{2\lambda}=\frac{3\|X\|}{2\lambda}.$ (46)
Note that $\frac{3\|X\|}{2\lambda}\leq\frac{1}{2}$ since $\lambda>3\sqrt{A}\|X\|$ and we can apply (45) to evaluate
$\|W\|$ $\leq\frac{\|(X/\lambda-Z)(I+Y)^{-1}\|^{2}}{1-\|(X/\lambda-Z)(I+Y)^{-1}\|}\leq\frac{A\|(X/\lambda-Z)\|^{2}}{1-\sqrt{A}\|X/\lambda-Z\|}=\sqrt{A}\|X/\lambda-Z\|h(\sqrt{A}\|X/\lambda-Z\|)$ (47)
$\leq\sqrt{A}\frac{3\|X\|}{2\lambda}h\left(\frac{3\sqrt{A}\|X\|}{2\lambda}\right)\leq\frac{3\sqrt{A}\|X\|}{2\lambda}h\left(1/2\right)=\frac{3\sqrt{A}\|X\|}{2\lambda}$ (48)
where we used $\|(X/\lambda-Z)(I+Y)^{-1}\|\leq\|X/\lambda-Z\|\|(I+Y)^{-1}\|\leq\sqrt{A}\|X/\lambda-Z\|$ (since $(I+Y)^{-1}$ has eigenvalues smaller than 1) and $h(1/2)=1$.
Starting again with (42), we then have
$\|M(I+X/\lambda,Y)-(I+Y)^{-1}\|_{\infty}$ $\leq\|M(I+X/\lambda,Y)-(I+Y)^{-1}\|$ (49)
$=\|(Z-X/\lambda)(I+Y)^{-1}+W\|\leq\frac{3\sqrt{A}\|X\|}{2\lambda}+\frac{3\sqrt{A}\|X\|}{2\lambda}$ (50)
$=\frac{3\sqrt{A}\|X\|}{\lambda}.$ (51)
Let $\mathrm{DiagDom}(Y)=\min_{(ab)}\left((I+Y)^{-1}_{aa}-(I+Y)^{-1}_{ab}\right)$ which is strictly positive since $(I+Y)^{-1}$ is strictly max-diagonally dominant [9]. We therefore have that, for any $\lambda>3\sqrt{A}\|X\|/\mathrm{DiagDom}(Y)$,
$\|M(I+X/\lambda,Y)-(I+Y)^{-1}\|_{\infty}\leq\mathrm{DiagDom}(Y)$
and therefore $M(I+X/\lambda,Y)$ is max-diagonally dominant, as expected. ∎
Appendix F Inverse of super-Laplacian
Proposition 10.
Let $\Delta$ be a super-Laplacian matrix, then for any nodes $a\neq b$, we have
$e_{a}^{T}\Delta^{-1}e_{ab}=(\Delta^{-1})_{aa}-(\Delta^{-1})_{ab}\geq 0.$
Proof.
We prove the result by interpreting the coefficients of $\Delta^{-1}$ as a probability of some sample path of a discrete-time Markov process on the alternatives. Let $D$ be the diagonal of $\Delta$, and $P$ the matrix defined by
$\Delta=D(I-P).$ (52)
The matrix $P$ is a row-sub-stochastic matrix. Explicitly,
$P_{aa}$ $=0,$ $P_{ab}$ $=\frac{|\Delta_{ab}|}{\Delta_{aa}},$ $\sum_{b\in\mathcal{A}}P_{ab}$ $<1.$ (53)
Let $\bullet$ be an extra symbol and $\mathcal{A}_{\bullet}=\mathcal{A}\sqcup\{\bullet\}$. Let $\kappa_{a}=\Delta_{aa}-\sum_{b\neq a}|\Delta_{ab}|>0$. We define a Markovian random walk on $\mathcal{A}_{\bullet}$ by the transition matrix $T$:
$T(b|a)$ $=P_{ab}=\frac{|\Delta_{ab}|}{\Delta_{aa}},$ $T(\bullet|a)$ $=1-\sum_{b\in\mathcal{A}}P_{ab}=\frac{\kappa_{a}}{\Delta_{aa}},$ $T(a|\bullet)$ $=0,$ $T(\bullet|\bullet)$ $=1.$ (54)
Intuitively, this Markovian process walks over the alternatives according to the weights $|\Delta_{ab}|$, and at each step has a non-zero probability to end in the cemetery $\bullet$.
Now,
$\Delta^{-1}_{ba}$ $=\big{(}(1-P)^{-1}D^{-1}\big{)}_{ba},$ $\Delta^{-1}_{ba}$ $=\sum_{n\geq 0}\big{(}P^{n}D^{-1}\big{)}_{ba},$ $\Delta^{-1}_{ba}$ $=\sum P_{ba_{1}}\dots P_{a_{n-1}a}\frac{1}{\Delta_{aa}}.$ (55)
Therefore,
$\Delta^{-1}_{ba}\kappa_{a}=\sum T(a_{1}|b)\dots T(a|a_{n-1})T(\bullet|a).$ (56)
In other words, $\Delta^{-1}_{ba}\kappa_{a}$ is the probability that $a$ is the last alternative to be visited before the random walk is killed, given that it started at $b$. Notice that the alternative $a$ may be visited multiple times in those paths. Actually, any path that starts at $b$ and visits $a$ before being killed can be decomposed as the gluing of a path that starts at $b$ and reaches $a$, followed by a path that starts at $a$ and eventually revisits $a$ as its last step before being killed. Therefore,
$\Delta^{-1}_{ba}\kappa_{a}\leq\Delta^{-1}_{aa}\kappa_{a}.$ (57)
Since $\kappa_{a}>0$, and since $\Delta$ is symmetric, we finally obtain
$\Delta^{-1}_{aa}\geq\Delta^{-1}_{ab}.$ (58)
Appendix G Proof of Theorem 2
Proof.
Fix an arbitrary $\lambda>0$, and let $\mu=s^{2}+\lambda$. There exists a permutation matrix $P$, and an integer partition $A_{1}+\dots+A_{k}=A$, such that
$X$ $\triangleq\begin{pmatrix}x^{T}&sI\end{pmatrix}\begin{pmatrix}x\\ sI\end{pmatrix}+\lambda I$ (59)
$=x^{T}x+\mu I$ (60)
$=P\cdot\Big{(}\mu I+\text{block\_diagonal}(J_{A_{1}},\dots,J_{A_{k}})\Big{)}\cdot P^{-1}$ (61)
$=P\cdot\text{block\_diagonal}(\mu I_{A_{1}}+J_{A_{1}},\dots,\mu I_{A_{k}}+J_{A_{k}})\cdot P^{-1}$ (62)
where every $J_{A_{i}}$ is a matrix of size $A_{i}\times A_{i}$ with all its entries set to $1$. Up to renaming the alternatives, we can assume, without loss of generality, that $P=I$.
We notice that the all-one matrix $J$, say of size $A$, satisfies $J^{2}=AJ$, and
$(\mu I+J)\frac{1}{\mu}\Big{(}I-\frac{1}{A+\mu}J\Big{)}=1$ (63)
Therefore, $X$ is invertible and
$X^{-1}=\text{block\_diagonal}\Big{(}\frac{1}{\mu}\big{(}I_{A_{1}}-\frac{1}{A_{1}+\mu}J_{A_{1}}\big{)},\dots,\frac{1}{\mu}\big{(}I_{A_{k}}-\frac{1}{A_{k}+\mu}J_{A_{k}}\big{)}\Big{)}$ (64)
Thus, $X^{-1}$ is super-Laplacian. This proves that $\begin{pmatrix}x&sI\end{pmatrix}^{T}$ is a diffusion embedding. The monotonicity of $\textsc{GBT}_{f,\sigma,x,L}$ follows from Theorem 1. ∎